OpenStudy (anonymous):

the normal line to the graph y=x^4 +1 at x=1 also intersects the graph at which of the following values of x?

OpenStudy (anonymous):

slope of line you get by finding the derivative and replacing x by 1 then take the negative reciprocal to find the slope of the normal k?

OpenStudy (anonymous):

ok so i get -1/4, but i dont understand what to do next. do i use a graphing calc?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[y'=4x^3\] at 1 get 4, negative reciprocal is \[-\frac{1}{4}\] so that his the slope of the normal line at x = 1, y = 2 to the equation of the normal line is \[y-2=-\frac{1}{4}(x-1)\] or \[y=-\frac{1}{4}(x-1)+2\]

OpenStudy (anonymous):

obviously the normal line intersects the graph at \[(1,2)\] the question is where else does it intersect.

OpenStudy (anonymous):

and for that one you are on your own because you have a fourth degree equation to solve, and it is not trivial

OpenStudy (phi):

sat gave you most of the answer. You could use a graphing calculator to finish the problem. You have two curves: \[ y= x^4 +1 \] \[ y= -\frac{1}{4}x + \frac{9}{4} \] you have to find where they intersect.